### v3

# Strong and noise-tolerant entanglement in dissipative optomechanics

Abstract
Macroscopic entanglement, as critical quantum resources in quantum information science, has been extensively studied in coherent optomechanics over the past decades. However, entanglement in dissipative optomechanics, where the cavity linewidth is dependent on the position of the mechanical resonator, remains unexplored. In this work, we study quantum entanglement in a dissipative optomechanics, realized by a Michelson–Sagnac interferometer and a movable membrane. Such dissipative optomechanics allows to switch coherent and dissipative optomechanical coupling at will. With reliable parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power with coherent (dissipative) coupling. Further, we show that quantum entanglement generated by dissipative coupling is significantly stronger and more noise-tolerant than that generated by coherent coupling. When both the coherent and dissipative couplings are simultaneously considered, entanglement is instead weakened due to quantum interference. Our result indicates that dissipative optomechanical coupling may be a promising path to engineer strong and noise-tolerant quantum entanglement.
I Introduction
Coherent optomechanics, where the cavity frequency is dependent on the position of the mechanical resonator, has garnered siginificant interest theoretically and experimentally over the past two decades Aspelmeyer et al. (2014). In such systems, various proposals have been put forward, including ground-state cooling Teufel et al. (2011); Gröblacher et al. (2009); Schliesser et al. (2008), bistability Dorsel et al. (1983); Gozzini et al. (1985) and tristability Xiong et al. (2016), optomechancally induced transparency Weis et al. (2010); Safavi-Naeini et al. (2011); Kronwald and Marquardt (2013), quantum squeezing Liao and Law (2011); Nunnenkamp et al. (2010); Lü et al. (2015); Aggarwal et al. (2020); Purdy et al. (2013); Luo et al. (2022), phase transition Wang et al. (2024); Jäger et al. (2019); Mann et al. (2018); Xu et al. (2017), microwave-optical conversion Guo et al. (2016); Han et al. (2021); Sahu et al. (2022); Andrews et al. (2014); Chen et al. (2023a); Xu et al. (2016), spin-optomechanical interface Xiong et al. (2021a); Chen et al. (2021); Peng et al. (2023); Xiong et al. (2023); Jäger et al. (2019), higher-order exceptional point Xiong et al. (2021b, 2022), and macroscopic quantum entanglement Vitali et al. (2007); Tian (2012, 2013); Wang and Clerk (2013); Chen et al. (2024, 2023b); Liu et al. (2024a); Lai et al. (2022); Jiao et al. (2020, 2022); Liu et al. (2024b); Shang and Li (2024); Ockeloen-Korppi et al. (2018). Recently, dissipative optomechanics was theoretically proposed in a superconducting microwave circuit Elste et al. (2009), where the cavity linewidth is a function of the mechanical displacement Elste et al. (2009), giving rise to a dissipative coupling between the mechanical and cavity modes. Later, the setup was extended to a Michelson-Sagnac interferometers (MSI) with a movable membrane Xuereb et al. (2011), where coherent or dissipative optomechanical coupling is allowed to switch on and off at will by tuning the position of the membrane or the reflectivity of the beam splitter. Experimentally, dissipative optomechanics have been demonstrated in MSI Sawadsky et al. (2015); Friedrich et al. (2011), microdisc resonator coupled to a nanomechanical waveguide Li et al. (2009), photonic crystal split-beam nanocavity Wu et al. (2014), graphene drums coupled to high-Q microsphere Cole et al. (2015), and levitated sphere trapped by dual-beam optical tweezers Kuang et al. (2023). Within dissipative optomechanics, diverse intriguing quantum phenomena, such as unconventional bistability Kyriienko et al. (2014), quantum squeezing Qu and Agarwal (2015); Huang and Chen (2020), enhanced ground-state cooling Huang and Chen (2018); Liu et al. (2023), quantum sensing Huang and Agarwal (2017), negative-damping instability Elste et al. (2009), self-sustained oscillations Huang et al. (2018), and phonon lasing Kuang et al. (2023); Zhang et al. (2022) have been investigated. However, macroscopic quantum entanglement has not been revealed yet with dissipative optomechanical coupling, although it has been intensively studied in coherent optomechanics.
In this work, we propose to investigate quantum entanglement in an experimentally feasible dissipative optomechanical system Sawadsky et al. (2015); Friedrich et al. (2011). The setup consists of a fixed perfect mirror and an effectively movable mirror, realized by a Michelson–Sagnac interferometer (MSI) with an integrated membrane. We begin by analyzing the steady-state mechanical displacement under both coherent and dissipative coupling. Notably, with coherent coupling, the mechanical displacement exhibits a nonlinear dependence on the driving power, consistent with conventional optomechanics Aspelmeyer et al. (2014). In contrast, under dissipative coupling, the displacement shows a linear dependence on the driving power. Furthermore, we explore quantum entanglement in the presence of coherent and dissipative coupling. Interestingly, we find that entanglement induced by dissipative coupling is not only significantly stronger than that generated by coherent coupling but also remarkably more robust against bath temperature. For entanglement with coherent optomechanical coupling, its survival temperature is approximately $\sim 8.5\,\text{K}$, while for entanglement with dissipative optomechanical coupling, it can be up to $\sim 25\,\text{K}$. When both the coherent and dissipative couplings are simultaneously considered, entanglement instead reduces due to quantum interference. Our work provides a promising path to generate strong and noise-resilient quantum entanglement in quantum systems with dissipative coupling.
The rest paper is organized as follows. In Sec. II, we introduce the proposed model and its Hamiltonian. Then, the steady-state solutions of the system are investigated in Sec. III. In Sec. IV, the dynamics of the fluctuation and covariance matrix are given. In Secs. V, quantum entanglement is investigated. Finally, a conclusion is given in Sec. VI.
II The model and Hamiltonian
Figure 1: (a) Schematic diagram of a typical dissipative optomechanical system, realized by a perfect fixed mirror $M_{1}$ and an effective movable mirror $\mathcal{M}$. (b) The structure of the effective movable mirror $\mathcal{M}$. The effective mirror consists of a Michelson-Sagnac interferometer (made up of a beam-splitter and two mirrors $M_{2}$ and $M_{3}$) and a movable membrane (mechanical resonator). $L$ is the effective length of the cavity ($2L$ is the length of the Sagnac mode ${\rm M_{1}-BS-M_{2}-M_{3}-BS}$). $a_{\rm in}$ is the input port and $a_{\rm out}$ is the output port.
As shown in Fig. 1(a), we consider an experimentally feasible dissipative optomechanics Sawadsky et al. (2015); Friedrich et al. (2011), consisting of a fixed mirror $M_{1}$ and an effective movable mirror $\mathcal{M}$. The effective mirror can be realized by a Michelson-Sagnac interferometer (MSI) and a movable membrane between mirrors $M_{2}$ and $M_{3}$ [see Fig. 1(b)]. $L$ is the effective length of the cavity ($2L$ is the length of the Sagnac mode ${\rm M}_{1}-{\rm BS}-{\rm M}_{2}-{\rm M}_{3}-{\rm BS}$). The membrane is treated quantum mechanically, with its displacement and momentum operators denoted by $x$ and $p$, respectively, satisfying the commutation relation $[x,p]=i\hbar$. Upon the introduction of an external driving field (with frequency $\omega_{d}$ and power $P$) at the beam splitter, the system can be effectively modeled as a typical dissipative optomechanical system. Thus, both the cavity frequency $\omega_{a}(x)$ and the decay rate $\kappa_{a}(x)$ depend on the displacement $x$ of the membrane. For small membrane displacements $x$, we can Taylor expand $\omega_{a}(x)$ and $\kappa_{a}(x)$ to the first order, i.e., $\omega(x)\approx\omega_{a}-g_{\omega}x$ and $\kappa(x)\approx\kappa_{a}-g_{\kappa}x$, where $\omega_{a}$ ($\kappa_{a}$) is the frequency (decay rate) of the cavity with annihilation (creation) operator $a$ ($a^{\dagger}$) when the membrane is at its equilibrium. The Hamiltonian of the considered system is given by Elste et al. (2009); Xuereb et al. (2011)
$H_{S}=$ $(\omega_{a}-g_{\omega}x)a^{\dagger}a+\frac{1}{2}\omega_{m}(x^{2}+p^{2})$
[…]
where $\omega_{m}$ is the frequency of the membrane, $\omega$ is the frequency of the bosonic mode (i.e., bath) with annihilation (creation) operator $a_{\omega}$ ($a_{\omega}^{\dagger}$), $x^{\prime}=\sqrt{2\kappa_{a}}-{g_{\kappa}x}/{\sqrt{2\kappa_{a}}}$ is the position-dependent coupling strength between the optical cavity and the bosonic mode. In fact, the bosonic mode here is to introduce a zero-value expectation input field $a_{\rm in}=\int\frac{d\omega}{\sqrt{2\pi}}e^{-i\omega(t-t_{0})}a_{\omega,0}$, with $a_{\omega,0}$ the operator at an initial time, to the cavity mode $a$, which can be easily obtained following the standard input-output theory with the Markov approximation. At the optical domain, the input field satisfies the relations Walls and Milburn (1994)
$\langle a_{\rm in}(t)a_{\rm in}(t^{\prime})\rangle=$ $0,$
$\langle a_{\rm in}(t)a^{\dagger}_{\rm in}(t^{\prime})\rangle=$ $\delta(t-t^{\prime}).$ (2)
Substituting the input field $a_{\rm in}$ into Eq. (II), the Hamiltonian of the driven unconventional MSI, with respect to the driving frequency $\omega_{d}$, becomes
$\mathcal{H}_{S}=$ $(\Delta_{a}-g_{\omega}x)a^{\dagger}a+\frac{1}{2}\omega_{m}(x^{2}+p^{2})$
$+ix^{\prime}(A_{\rm in}a^{\dagger}-A_{\rm in}^{\dagger}a),$ (3)
where $\Delta_{a}=\omega_{a}-\omega_{d}$ is the frequency detuning of the cavity mode from the driving field, and
$A_{\rm in}=a_{\rm in}+\mathcal{E}.$ (4)
Here $\mathcal{E}=|\mathcal{E}|\exp(i\theta)$, with $|\mathcal{E}|=\sqrt{\mathcal{P}/\hbar\omega_{d}}$ being the amplitude of the driving field and $\theta$ the phase. Notably, $\mathcal{E}$ can also be interpreted as the steady-state value of the operator $A_{\rm in}$, while $a_{\rm in}$ represents the fluctuation of $A_{\rm in}$.
III Steady-state solution
When dissipation is included, the dynamics of the proposed system can be governed by the quantum Langevin equantion $\dot{\mathcal{O}}=-i[\mathcal{O},\mathcal{H}_{S}]+{Dissipation}+{Noise}$, specifically,
$\dot{x}=$ $\omega_{m}p,$
$\dot{p}=$ $g_{\omega}a^{\dagger}a-\omega_{m}x+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}(A_{\rm in}^{\dagger}a-A_{\rm in}a^{\dagger})-\gamma_{m}p+\xi,$
$\dot{a}=$ $-(\kappa_{a}+i\Delta_{a})a+ig_{\omega}ax+g_{\kappa}ax+x^{\prime}A_{\rm in},$ (5)
where $\gamma_{m}$ is the damping rate of the membrane and $\xi$ is the Brownian noise operator arising from the coupling of the membrane oscillator to the hot environment. The correlation function of $\xi$ is given by Giovannetti and Vitali (2001) $\langle\xi(t)\xi(t^{\prime})\rangle=\frac{\gamma_{m}}{\omega_{m}}\int\frac{d\omega}{2\pi}e^{-i\omega(t-t^{\prime})}\omega[{\rm coth}(\frac{\hbar\omega}{2k_{b}T})+1]$. In the limit of the high quality factor, $\omega_{m}\gg\gamma_{m}$, the Brownian noise $\xi$ becomes am approximate Markovian noise Benguria and Kac (1981), i.e.,
$\langle\xi(t)\xi(t^{\prime})+\xi(t^{\prime})\xi(t)\rangle/2\backsimeq\gamma_{m}(2n_{\rm th}+1)\delta(t-t^{\prime})$ (6)
where the mean thermal phonon number $n_{\rm th}$ = $[{\rm exp}(\hbar\omega_{m}/{k_{B}T}-1)]^{-1}$ with $k_{B}$ the Boltzmann constant and $T$ the bath temperature.
Figure 2: The normalized steady-state displacment $x_{s}$ of the membrane vs (a) the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ and (b) the power of the driving field. The red curve corresponds to the case of the coherent coupling ($g_{\omega}/2\pi=2$ Hz). The blue curve is plotted for the dissipative coupling ($g_{\kappa}/2\pi=2$ Hz). In (a) and (b), $\Delta_{a}=3\kappa_{a}$ is assumed.
When the evolution time is long enough, the system reaches its steady state, denoted by $a_{s}$, $x_{s}$, and $p_{s}$. From Eq. (III), the steady-state equation can be directly obtained as
$\dot{x}_{s}=$ $\omega_{m}p_{s},$
$\dot{p}_{s}=$ $g_{\omega}|a_{s}|^{2}-\omega_{m}x_{s}+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}(\mathcal{E}^{*}a_{s}-\mathcal{E}a_{s}^{*})-\gamma_{m}p_{s},$
$\dot{a}_{s}=$ $-(\kappa_{s}+i{\Delta}_{s})a_{s}+x_{s}^{\prime}\mathcal{E},$ (7)
where ${\Delta}_{s}=\Delta_{a}-{g_{\omega}}x_{s}$, $\kappa_{s}=\kappa_{a}-g_{\kappa}x_{s}$, and $x_{s}^{\prime}=\sqrt{2\kappa_{a}}-{g_{\kappa}x_{s}}/{\sqrt{2\kappa_{a}}}$. At the steady state, we have $\dot{\mathcal{O}}_{s}=0$, equivalently
$\omega_{m}p_{s}=$ $0,$
$g_{\omega}|a_{s}|^{2}-\omega_{m}x_{s}+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}(\mathcal{E}^{*}a_{s}-\mathcal{E}a_{s}^{*})-\gamma_{m}p_{s}=$ $0,$ (8)
$-(\kappa_{s}+i{\Delta}_{s})a_{s}+x_{s}^{\prime}\mathcal{E}=$ $0$
By solving these equations, we have
$p_{s}=$ $0,~{}~{}a_{s}=\frac{x_{s}^{\prime}\mathcal{E}}{\kappa_{s}+i{\Delta}_{s}},$
$x_{s}=$ $\frac{g_{\omega}n_{s}}{\omega_{m}}+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}\cdot\frac{a_{s}\mathcal{E}^{*}-a_{s}^{*}\mathcal{E}}{\omega_{m}}.$ (9)
Obivously, the steady-state displacement of the membrane is determined by both the dispersive and dissipative coupling between photons and phonons. For a coherent coupling ($g_{\kappa}=0$), $x_{s}$ is proportional to the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ [see the red curve in Fig. 2(a)]. For a dissipative coupling ($g_{\omega}=0$), $x_{s}$ becomes
$x_{s}=[1-\sqrt{1-2(g_{\kappa}/\kappa_{a})^{2}(\Delta_{a}/\omega_{m})n_{s}}]\omega_{m}\kappa_{a},$ (10)
where $x_{s}\neq 2\kappa_{a}/g_{\kappa}$ and $n_{s}\leq(\kappa_{a}/g_{\kappa})^{2}(\omega_{m}/2\Delta_{a})$ are both required. When $n_{s}=0$, $x_{s}=0$, i.e., the membrane is positioned at its equilibrum. Here $x_{s}$ and $n_{s}$ are quadratic dependence for the dissipative coupling, demonstrated by the blue curve in Fig. 2(a). One may further notice that the displacement of the membrane caused by the coherent coupling is larger than that caused by the dissipative coupling, as clearly shown in Fig. 2(a). To plot Fig. 2(a), we adopt typical experimental parameters Sawadsky et al. (2015): the wavelength of the driving field $\lambda=1064$ nm, $\omega_{d}/2\pi=c/\lambda=281.96$ THz, the quality factor of the membrane $\mathcal{Q}_{m}=5.8\times 10^{5}$, $\omega_{m}/2\pi=136$ kHz, $\gamma_{m}/2\pi=0.23$ Hz, $L=8.7$ cm, $m=80$ ng, and $\kappa_{a}/2\pi=1.5$ MHz. These experimental parameters ensure the stability of the proposed system according to Routh-Hurwitz criterion.
From Eq. (III), one can see that $x_{s}$ actually satisfies the following cubic equation related to the amplitude of the driving field $\mathcal{E}$,
$ax_{s}^{3}+bx_{s}^{2}+cx_{s}+d=0,$ (11)
where
$a=$ $2(g_{\omega}^{2}+g_{\kappa}^{2})\omega_{m}\kappa_{a},$
$b=$ $-4\omega_{m}\kappa_{a}(g_{\omega}\Delta_{a}+g_{\kappa}\kappa_{a})-3g_{\omega}g_{\kappa}^{2}|\mathcal{E}|^{2},$
$c=$ $2\omega_{m}\kappa_{a}(\kappa_{a}^{2}+\Delta_{a}^{2})+2g_{\kappa}|\mathcal{E}|^{2}(g_{\kappa}\Delta_{a}+4g_{\omega}\kappa_{a}),$ (12)
$d=$ $-4\kappa_{a}|\mathcal{E}|^{2}(g_{\omega}\kappa_{a}+g_{\kappa}\Delta_{a}).$
Explicitly, Eq. (11) may have three real roots, corresponding to bistability. But in fact, the bistability can only be predicted for coherent coupling, as shown by the red curve in Fig. 2(b). This indicates that dissipative coupling can not generate the typical bistability in optomechanics, demonstrated by the blue curve in Fig. 2(b).
IV Covariance matrix
With steady-state values in hand, we further expand each operator $\mathcal{O}$ around its steady state, i.e., $\mathcal{O}=\delta\mathcal{O}+\mathcal{O}_{s}$. Substituting this expression back into Eq. (III), we then obtain a set of equations for fluctuations ($\delta\mathcal{O}$). For the strong driving field ($n_{s}\gg 1$), the higher-order fluctuation terms can be safely neglected Vitali et al. (2007), corresponding to linearization. As a result, the dynamics for fluctuations can be written as
$\dot{\delta x}=$ $\omega_{m}\delta p,$
$\dot{\delta p}=$ $\frac{G_{\omega}\delta a^{\dagger}+G_{\omega}^{*}\delta a}{\sqrt{2}}-\omega_{m}\delta x-\gamma_{m}\delta p+\xi$
$+\frac{ig_{\kappa}(\mathcal{E}^{*}\delta a-\mathcal{E}\delta a^{\dagger})}{\sqrt{2\kappa_{a}}}+\frac{i(G_{\kappa}\delta a_{\rm in}^{\dagger}-G_{\kappa}^{*}\delta a_{\rm in})}{2\sqrt{\kappa_{a}}},$ (13)
$\dot{\delta a}=$ $-(\kappa_{s}+i\Delta_{s})\delta a+(\frac{G_{\kappa}+iG_{\omega}}{\sqrt{2}}-\frac{g_{\kappa}\mathcal{E}}{\sqrt{2\kappa_{a}}})\delta x+x_{s}^{\prime}a_{\rm in},$
where $G_{\omega(\kappa)}=\sqrt{2}g_{\omega(\kappa)}a_{s}$ is the linearized dispersive (dissipative) coupling strength, enhanced by the amplitude of the driving field. By further defining
$\delta x_{a}=$ $\frac{\delta a^{\dagger}+\delta a}{\sqrt{2}},~{}~{}\delta p_{a}=i\frac{\delta a^{\dagger}-\delta a}{\sqrt{2}},$
$x_{\rm in}=$ $\frac{a_{\rm in}^{\dagger}+a_{\rm in}}{\sqrt{2}},~{}~{}p_{\rm in}=i\frac{a_{\rm in}^{\dagger}-a_{\rm in}}{\sqrt{2}},$ (14)
Eq. (IV) can be rewritten in a more compact form, i.e.,
$\dot{u}(t)=Au(t)+n(t),$ (15)
where $u^{T}(t)=(\delta x,\delta p,\delta x_{a},\delta p_{a})$ is the vector of fluctuation operators, $n(t)$ is the vector of noise operator, given by $n^{T}(t)=(0,\xi+\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}p_{\rm in},x_{s}^{\prime}x_{\rm in},x_{s}^{\prime}p_{\rm in})$, and the drift matrix $A$ is
$A=\left(\begin{array}[]{cccc}0&\omega_{m}&0&0\\ -\omega_{m}&-\gamma_{m}&G_{\omega}+\frac{G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&-\frac{G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}\\ G_{\kappa}-\frac{2G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&0&-\kappa_{s}&\Delta_{s}\\ G_{\omega}-\frac{2G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&0&-\Delta_{s}&-\kappa_{s}\\ \end{array}\right).$ (16)
To obtain the drift matrix $A$, we have taken $a_{s}$ as real number, which can be realized by setting $\tan\theta=\Delta_{a}/\kappa_{a}$ in Eq. (III).
Since the phonon noise $\xi$ and the photon noise $a_{\rm in}$ are zero-mean quantum Gaussian noises and the dynamics is linearized [see Eq. (IV)], the quantum steady state for fluctuations is a zero-mean bipartite Gaussian state, fully described by a $4\times 4$ correlation matrix ${V}$, where $V_{ij}=\langle u_{i}(t)u_{j}(t^{\prime})+u_{j}(t^{\prime})u_{i}(t)\rangle/2$ with $i,j=1,2,3,4$. The $V$ is obtained by solving the Lyapunov equation $AV+VA^{T}=-D$, where $D$ is the diffusion matrix, defined by $\langle n_{i}(t)n_{j}(t^{\prime})+n_{j}(t^{\prime})n_{i}(t)\rangle/2=D_{ij}\delta(t-t^{\prime})$. More specifically, $D$ can be expressed as
$D=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&\gamma_{m}(2n_{\rm th}+1)+\frac{G_{\kappa}^{2}}{4\kappa_{a}}&0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{x_{s}^{\prime}}{2}\\ 0&0&\frac{x_{s}^{\prime,2}}{2}&0\\ 0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{x_{s}^{\prime}}{2}&0&\frac{x_{s}^{\prime,2}}{2}\\ \end{array}\right).$ (17)
Compared to coherent coupling case, dissipative coupling strength appears in the off-diagonal elements of the diffusion matrix, playing an crucial role in generation of bipartite entanglement below.
To measure optomechanical entanglement, the logarithmic negativity ($E_{N}$) is taken Adesso et al. (2004), i.e.,
$E_{N}\equiv\rm max[0,-ln2\eta^{-}],$ (18)
where $\eta^{-}=2^{-1/2}[\Sigma({V})-(\Sigma({V})^{2}-4\rm det{{V}})^{1/2}]^{1/2}$ with $\Sigma({V})=\rm{det}{A}+\rm{det}B-2\rm{det}C$, ${{V}}=\begin{pmatrix}A&C\\ {C^{T}}&B\\ \end{pmatrix}$ is the $2\times 2$ block form. When $E_{N}>0~{}(=0)$, the state is entangled (separable). Moreover, the larger $E_{N}$, the stronger entanglement.
V Optomechanical Entanglement
To investigate optomechanical entanglement, we consider three specific scenarios: (i) coherent coupling, (ii) dissipative coupling, and (iii) cooperative coupling (i.e., both the coherent and dissipative couplings are considered).
V.1 Coherent coupling ($g_{\kappa}=0$)
In the coherent coupling case ($g_{\kappa}=0$), we present the entanglement measure $E_{N}$ as a function of both the coherent coupling strength $g_{\omega}$ and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$ in Fig. 3(a) to explore the conditions for optimal optomechanical entanglement at a bath temperature of $T=0.4$ K. The results reveal that the generated entanglement remains relatively weak compared to that typically observed in the conventional optomechanical system Vitali et al. (2007). This is because the stronger coupling strength will lead to system unstable. To numerically ensure system stable, the coupling strength is required to be weak, giving rise to weak entanglement. Nonetheless, optimal entanglement can still be achieved around the point ($\Delta_{s}/\omega_{m},g_{\omega}/2\pi)\approx(2,3.1)$, as marked by the green star, indicating that even under these constraints, entanglement can be optimized.
Figure 3: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the coherent coupling strength $g_{\omega}$ at bath temperature $T=400$ mK. (b) The logarithmic negativity $E_{N}$ vs the coherent coupling strength $g_{\omega}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=2$ is fixed. (c) The logarithmic negativity $E_{N}$ vs the bath temperature $T$ with $g_{\omega}/2\pi=3.1$ Hz and $\Delta_{s}/\omega_{m}=2$. Other parameters are the same as those in Fig. 2 except for $\mathcal{P}=50$ mW.
To further investigate the coherent coupling effect on entanglement, we plot $E_{N}$ as a function of $g_{\omega}$ in Fig. 3(b), where $\Delta_{s}\approx 2\omega_{m}$ is fixed. It is obviously shown that $E_{N}$ increases with $g_{\omega}$. Once up to its maximum, $E_{N}$ will have a sharp reduction by further increasing $g_{\omega}$. Such abrupt behavior corresponds to the transition from the stability to instability, as shown by the pink-shaded region. Additionally, when the bath temperature increases to $T=4$ K [see the blue dashed curve], the coupling for generation of quantum entanglement is required to be larger and $E_{N}$ has a significant reduction. This indicates that quantum entanglement can be affected by the bath temperature. To clearly evaluate the effect of bath temperature, we plot $E_{N}$ as a function of the bath temperature in Fig. 3(c), where $g_{\omega}/2\pi=3.1$ Hz is fixed to obtain optimal entanglement. As the temperature increases, $E_{N}$ gradually declines. When the bath temperature increases to $T\approx 8.5$ K, $E_{N}=0$, that is, quantum entanglment vanishes and the system becomes separable.
V.2 Dissipative coupling ($g_{\omega}=0$)
In dissipative coupling case ($g_{\omega}=0$), we observe that the optimal entanglement can be significantly stronger than in the coherent coupling case, even surpassing the entanglement achieved in conventional optomechanical system (i.e., $E_{N}=0.35$Vitali et al. (2007). This is demonstrated in Fig. 4(a), where $E_{N}$ is plotted as a function of both the dissipative coupling strength ($g_{\kappa}$) and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$. The optimal entanglement is found at approximately ($\Delta_{s}/\omega_{m},g_{\kappa}/2\pi)\approx(0.1,20)$, as indicated by the green star. In contrast to the coherent coupling case [see Fig. 2(a)], where the optimal entanglement occurs near the mechanical red sideband at $\Delta_{s}/\omega_{m}=2$, the optimal entanglement in the dissipative coupling case is obtained when the cavity is nearly resonant with the driving field, specifically at $\Delta_{s}/\omega_{m}=0.1$, which is far removed from the mechanical red sideband.
Figure 4: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the dissipative coupling strength $g_{\kappa}$ at bath temperature $T=400$ mK. (b) The logarithmic negativity $E_{N}$ vs the dissipative coupling strength $g_{\kappa}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=0.1$ is fixed. (c) The logarithmic negativity $E_{N}$ vs the bath temperature $T$ with $g_{\kappa}/2\pi=19$ Hz and $\Delta_{s}/\omega_{m}=0.1$. Other parameters are the same as those in Fig. 2 except for $\mathcal{P}=50$ mW.
Furthermore, in Fig. 4(b), we investigate the behavior of the optomechanical entanglement as a function of the dissipative coupling strength, while fixing $\Delta_{s}=0.1\omega_{m}$. The resulting behavior of $E_{N}$ mirrors that observed in the coherent coupling case (see the red curve). The abrupt drop in entanglement marks the transition of the system from a stable to an unstable region, as indicated by the pink area. Notably, a relatively large dissipative coupling strength is required to achieve the optimal entanglement. This suggests that in the regime dominated by dissipative coupling, stronger interactions are necessary to generate significant quantum effects. As the bath temperature increases, for instance to $T=4$ K, the entanglement is reduced (see the blue curve). To explore the influence of bath temperature on optomechanical entanglement, we fix $g_{\kappa}/2\pi=19$ Hz and present the results in Fig. 4(c). We observe a sharp decrease in entanglement as the system temperature increases from 0 K to 5 K. However, beyond this point, the entanglement becomes more tolerant to further temperature increases, suggesting that the dissipative coupling becomes more robust to thermal noise at higher temperatures. Remarkably, entanglement persists even up to 25 K, This indicates that strong, noise-resistant entanglement can be generated through dissipative optomechanical coupling, opening new possibilities for quantum information processing and communication in noisy
V.3 Cooperative coupling ($g_{\omega}\neq 0$ and $g_{\kappa}\neq 0$)
Figure 5: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the ratio of $g_{\kappa}/g_{\omega}$ at bath temperature $T=400$ mK. (b) The logarithmic negativity $E_{N}$ vs the ratio of $g_{\kappa}/g_{\omega}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=2.2$ is fixed. (c) The logarithmic negativity $E_{N}$ vs the bath temperature $T$ with $g_{\kappa}/g_{\omega}=0.01$ (see the red curve) and $g_{\kappa}/g_{\omega}=0.1$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=2.2$. Other parameters are the same as those in Fig. 2 except for $\mathcal{P}=50$ mW, $g_{\omega}/2\pi=3$ Hz.
To examine the effect of the cooperative coupling (i.e., coherent and dissipative couplings) on optomechanical entanglement, we set the coherent coupling strength to be $g_{\omega}/2\pi=3$ Hz, corresponding to the optimal entanglement condition at $\Delta_{s}/\omega_{m}=2$ in the dispersive coupling case [refer to Fig. 3(a)]. By progressively increasing the dissipative coupling strength, we observe a gradual decline in the entanglement, as shown in Fig. 5(a). The reduction in entanglement possibly arises from the interference between the coherent and dissipative couplings. When the bath temperature increases to $T=4$ K, it is observed that the optomechancial entanglement is further suppressed, as depicted by the blue dashed curve in Fig. 5(b). Moreover, the larger $g_{\kappa}$ is not benificial to entanglement generation. This can also be revealed by Fig. 5(c), where $g_{\kappa}/g_{\omega}=0.01$ (the red curve) and $g_{\kappa}/g_{\omega}=0.1$ (the blue curve) are taken . For $g_{\kappa}/g_{\omega}=0.1$, the survival temperature of the entanglement is about $5$ K. But for $g_{\kappa}/g_{\omega}=0.01$, the survival temperature can be up to $8$ K.
VI Conclusion
In a summary, we investigate quantum entanglement in a dissipative optomechanical system, realized by a fixed mirror and a Michelson–Sagnac interferometer with a membrane. This system allows the dissipative and coherent couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter, enabling us to explore quantum entanglement in three distinct scenarios: coherent coupling case, dissipative case, and cooperative case. Our results reveals that, in the coherent coupling case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, whereas in the dissipative coupling case, the displacement shows a linear dependence. Furthermore, entanglement generated through dissipative coupling is not only significantly stronger but also more noise-tolerant compared to that achieved via coherent coupling. When both coupling mechanisms coexist, quantum entanglement is reduced, likely due to the interference between the two coupling interactions. These findings suggest that dissipative coupling may be a promising mechanism to engineer strong and noise-tolerant quantum entanglement.
This work is supported by the National Natural Science Foundation of China (Grant No. 12175001 ), the “Pioneer” and “Leading Goose” R&\&&D Program of Zhejiang (Grant No. 2025C01028), the Natural Science Foundation of Zhejiang Province (Grant No. LY24A040004), and the Shenzhen International Quantum Academy (Grant No. SIQA2024KFKT010).